# Neural Honeytrace: A Robust Plug-and-Play Watermarking Framework against Model Extraction Attacks

### v3

@@ 4.3 Multi-step Watermark Transmission
Training-free watermark embedding enables plug-and-play watermarking, however, according to Q2-A2 and Q3-A3, the channel capacity is not always ideal for watermark transmission. Therefore, we propose the multi-step watermark transmission strategy to enhance the robustness of Neural Honeytrace.
@@
where is used to balance model availability and watermark success rate, Bernoulli randomly samples with probability to decide whether to flip the label, and is a small random value to maintain randomness. Intuitively, for , the input sample will be labeled as the target class with a high probability. And by controlling , defenders can determine the flipping ratio for samples with . Eq. 10 links the predicted labels to the watermark similarities. As a results, Neural Honeytrace transmits watermark information to the label distribution of the surrogate dataset owned by attackers.
Robustness analysis. Label-flipping-based watermark information transmission embeds similarity scores in the label distribution of a set of samples with similar similarity scores. Consider the queries of these samples as a one-time message transmission, the channel capacity satisfies . And in practice, is usually a large number because attackers will utilize many samples with similar similarity scores to query the target model. Consequently, enables successfully watermark transmission under hard-label scenarios.
As for adaptive attacks, the variance of the mean of the noise in multi-step transmissions satisfies , following Eq. 5 and Eq. 6, the decreased noise variance will strictly increase the channel capacity, resulting in robust watermark transmission against adaptive MEAs.
Combining training-free watermark embedding and multi-step watermark transmission, Algorithm 1 summarizes the workflow of Neural Honeytrace.
@@
5 Experiments
In this section, we evaluate Neural Honeytrace against different model extractions and compare it with previous watermarking methods. We first introduce the experiment setup in Sec. 5.1. Then we compare the overall performance of different watermarking strategies against adaptive attacks in Sec. 5.2. In Sec. 5.3, we provide ablation study to further analyze Neural Honeytrace.
5.1 Experiment Setup
Datasets. We use four different image classification datasets to train the target model: CIFAR-10 [24], CIFAR-100 [24], Caltech-256 [25], and CUB-200 [26].
We use another two datasets as surrogate datasets used by model extractions attackers, TinyImageNet-200 [27] for querying target models trained on CIFAR-10 and CIFAR-100, ImageNet-1K [28] for querying target models trained on Caltech-256 and CUB-200.
Models. We use two model architectures to train target models: VGG16-BN [29] for CIFAR-10 and CIFAR-100, and ResNet50 [30] for Caltech-256 and CUB-200. Following previous model extraction defenses [31, 11], the same architectures are used for surrogate models.
Metrics. We use three metrics to evaluate the effectiveness of different watermarking methods:
1.
Protected Accuracy indicates the accuracy of the protected model on clean samples, i.e., clean accuracy.
2.
Extracted Accuracy (E-Acc) indicates the clean accuracy of the stolen model reconstructed by attackers.
3.
Watermark Success Rate (WSR) measures the proportion of successfully activated watermarks. Specifically, it is the ratio of samples with added triggers that are classified into the target classes by the watermarked model but classified into different classes by non-protected models.
@@ 5.1 Experiment Setup
| Query | Attack | DAWN | EWE | Composite | MEA-Defender | Ours | |||||
| E-Acc | WSR | E-Acc | WSR | E-Acc | WSR | E-Acc | WSR | E-Acc | WSR | ||
| Protected Accuracy | 81.95% | 82.40% | 82.41% | 82.67% | 80.13% | ||||||
Model extraction attack methods. We consider two different query strategies, KnockoffNet [4] and JBDA-TR [5], as basic model extraction attack methods. KnockoffNet utilizes natural samples to query the target model and construct surrogate training datasets, while JBDA-TR uses Jacobian-based data augmentation to generate synthetic data and probe the decision boundary of the target model.
Additionally, we consider the following adaptive attack methods on the basis of KnockoffNet and JBDA-TR:
1.
S4L Attack [7]: The loss function consists a CE loss and a semi-supervised loss, which helps train the model on both labeled and unlabeled data.
2.
Smoothing Attack [9]: Each sample is augmented and fed into the target model times, and the prediction is computed as the average of queries. In experiments of this paper, we set .
3.
D-DAE [10]: Attackers train a defense detection model and a label recover model to detect and bypass potential defenses. In the default configuration, we use recover models trained on different output perturbation methods. For advanced attacks (oracle attack in Sec. 5.3), we train the recover model on different watermarking strategies.
4.
p-Bayes Attack [11]: Attackers use independent and neighborhood sampling to perform Bayes-based estimation for original labels.
5.
Top-1 Attack: Only hard-labels are used for training surrogate models, which can be also considered as a defense mechanism if defenders only provide hard-labels.
Watermarking strategies. We compare Neural Honeytrace with the following watermarking strategies in experiments:
1.
DAWN [17]: Defenders randomly flip the predicted label of a subset of input queries and record these sample-label pairs as watermarks. In the default configuration, we assume that of all queries are executed by benign users and by model extraction attackers.
2.
EWE [16]: Defenders utilize the Soft Nearest Neighbor Loss (SNNL) to minimize the distance between watermark features and natural features.
3.
Composite Backdoor [18]: Spliced in-distribution samples are used as watermarks (triggers) to increase the watermark success rate.
4.
MEA-Defender [20]: Defenders introduce the utility loss, the watermarking loss, and the evasion loss to balance the model availability and watermark success rate.
5.2 Experimental Results
Overall performance comparison. We begin by comparing Neural Honeytrace with four baseline methods across four different datasets and six model extraction attack strategies. The experimental results for target models trained on CIFAR-10, CIFAR-100, Caltech-256, and CUB-200 are presented in Tab. II, Tab. III, Tab. IV, and Tab. V, respectively. For each defense, we report both the average and maximum extraction accuracy, as well as the average and minimum Watermark Success Rate (WSR). This approach highlights the effectiveness of various watermarking strategies under both optimal and worst-case scenarios, simulating an average attacker and more sophisticated attackers who select the most effective attack methods. We make the following additional observations based on the experiment results:
Existing watermarking methods are sensitive to dataset scale and data complexity. As shown in Tab. II and Tab. III, for smaller datasets and simpler data, existing watermarking methods are effective against Naive Attacks. However, for larger and more complex datasets, as demonstrated in Tab. IV, in four baseline methods, only MEA-Defender maintains acceptable success rates against Naive Attack. This is due to the increased complexity of the data, which leads to an increase in the information entropy of the output prediction. As a result, the watermark transmission process is subject to more channel noise and requires more robust watermark embedding strategies. Intriguingly, sometimes the sensitivity of dataset may also lead to higher watermark success rate. As shown in Tab. V, the watermark success rate of Composite Backdoor and MEA-Defender against Top-1 Attack is much higher than other methods, while the extraction accuracy decreases significantly. This is because the features of watermarks and query samples are similar in target models protected by these two defenses, resulting in mislabeling a large mount of model extraction queries during attacks.
Existing watermarking methods are vulnerable to adaptive attacks. The experimental results show that five adaptive attack strategies can weaken the watermark to varying degrees. Overall, the Smoothing Attack and Top-1 Attack have the most significant impact on watermark success when KnockoffNet and JBDA-TR are used as query strategies, respectively. For example, in Tab. II, for KnockoffNet+Smoothing Attack and JBDA-TR+Top-1 Attack, the average WSR of four baseline methods is and , respectively. According to Eq. 2, a decrease in WSR leads to an exponential increase in the sample size required for an ownership claim. Therefore, existing methods are highly likely to fail in detecting stolen models if attackers have prior knowledge of potential defenses.
MEA-Defender outperforms other baseline defenses. Among four baseline methods, although DAWN achieves the highest average WSR across four datasets, it relies on the assumption that model extraction queries occur in a high percentage of all queries. However, in real world scenarios, the proportion of clean queries should be significantly larger than the proportion of extraction queries as the time the model is online grows, in which case the WSR of DAWN will decrease over time. Therefore, it cannot be used as a stable watermarking method. Compared with EWE and Composite Backdoor, MEA-Defender shows higher a average WSR and better robustness against most adaptive attacks. However, the minimum WSR for MEA-Defender indicates that it will still fail under certain adaptive attacks.
@@ 5.2 Experimental Results
Neural Honeytrace outperforms existing defenses. Compared to existing watermarking strategies, Neural Honeytrace achieves better transferability across different datasets. As shown by comparing Tab. II and Tab. III, the watermark success rate of Neural Honeytrace remains stable as the number of output classes increases. For datasets with more complex features, such as those in Tab.IV and Tab.V, Neural Honeytrace maintains its effectiveness, whereas most previous methods suffer from significant performance degradation. Meanwhile, Neural Honeytrace shows better robustness against different adaptive attacks, as reflected in the minimum watermarking success rate. Fig. 6 provides a visualization of the sample sizes required by different watermarking strategies for ownership claim. Considering both the average and worst case scenarios, it can be seen that Neural Honeytrace requires fewer queries compared to previous methods. For example, in the worst-case scenario, MEA-Defender and Neural Honeytrace require 193,252 and 1,857 queries respectively to achieve copyright declaration.
| Acc | WSR | Acc | WSR | Acc | WSR | Acc | WSR | Acc | WSR | Acc | WSR | ||
| Attack | Setting | CIFAR-10 | CIFAR-100 | ||
| Acc | WSR | Acc | WSR | ||
| DFME | Soft-label | 66.53% | 53.40% | 38.29% | 49.80% |
| DFMS | Soft-label | 71.81% | 62.00% | 40.78% | 57.40% |
| Hard-label | 68.69% | 68.20% | 36.45% | 66.60% | |
| DisGUIDE | Soft-label | 76.48% | 57.80% | 45.58% | 64.20% |
| Hard-label | 73.10% | 64.20% | 40.44% | 70.80% | |
| Attack | V-16R-18 | V-16R-50 | R-18V-16 | |||
| Acc | WSR | Acc | WSR | Acc | WSR | |
| Naive | 54.04% | 47.60% | 56.44% | 55.80% | 65.67% | 77.80% |
| S4L | 52.50% | 55.00% | 53.46% | 61.40% | 62.29% | 80.80% |
| Smoothing | 53.74% | 58.20% | 56.68% | 59.20% | 65.03% | 73.40% |
| DDAE | 41.93% | 54.80% | 44.76% | 56.20% | 65.59% | 68.20% |
| P-Bayes | 52.66% | 43.80% | 55.18% | 49.40% | 64.76% | 70.80% |
| Top-1 | 46.10% | 45.00% | 50.03% | 51.80% | 55.31% | 65.60% |
5.3 Ablation Study
In this section, we comprehensively evaluate the performance of Neural Honeytrace with different defense settings and against different attack settings.
Different Watermark Triggers. We first compare the watermark success rate of Neural Honeytrace using different watermark triggers. As discussed in Sec.4.2, Neural Honeytrace supports the transmission of various watermarks, and the form of these watermarks influences the cost of the watermarking process. In Tab.VI, we evaluate three different watermark triggers: white pixel blocks (used in EWE [16]), a semantic object (e.g., a specific copyright logo), and a composite trigger (the default configuration). For the first two trigger types, we adjust the hyperparameters to ensure that the protected model maintains accuracy similar to that of the default configuration.
As shown in Tab.VI, the composite trigger achieves the highest average watermark success rate among the three triggers, which aligns with the analysis in Sec.4.2. Compared to the semantic object trigger, white pixel blocks yield a higher average watermark success rate due to their simpler features, making them easier to learn. However, even the semantic object trigger remains robust against various adaptive attacks. Given its interpretable semantics (e.g., the owner’s logo), it could still have potential applications in real-world scenarios.
Different Query Datasets. Depending on the capabilities and prior knowledge, attackers may use different surrogate datasets to query the target model. In Tab.VI, we compare the performance of Neural Honeytrace on a target model trained on CIFAR-10 when attackers use CIFAR-10, CIFAR-100, and TinyImageNet as surrogate datasets, respectively. The experimental results show that for out-of-distribution surrogate datasets (CIFAR-100 and TinyImageNet), attackers achieve similar extraction accuracy, and Neural Honeytrace maintains high watermark success rates. However, when using the same training dataset as the surrogate (CIFAR-10), attackers achieve higher extraction accuracy, while the watermark success rate of Neural Honeytrace decreases. This occurs because Eq.8, which is used to detect potential extraction queries, fails in this case. Nevertheless, even in such scenarios, the watermark success rate remains sufficiently high for ownership claims within queries. And in real-world scenarios, it is less possible for attackers to get access to the training dataset of the target model.
Data-free Model Extraction Attacks. Data-free model extraction attacks generate synthetic data using a trained generator guided by gradient approximations to replicate a victim model’s behavior without access to real training data, making MEAs more achievable. Tab. VII lists the defensive performance of Neural Honeytrace against three different data-free MEAs, DFME [6], DFMS [32], and DisGUIDE [33]. The experimental results show that Neural Honeytrace is effective for defending all three kinds of data-free MEAs, with the average WSR of . This is because Neural Honeytrace does not rely on prior knowledge about the query dataset of attackers and is data-irrelevant.
Different Model Architectures. The architecture of the surrogate model used by attackers may be different from the architecture of the target model. Tab. VIII lists the performance of Neural Honeytrace on different teacher-student model pairs trained on CIFAR-10. Combining Tab. VIII and Tab. II, it can be observed that when the model structures are different, the stolen accuracy will slightly decrease, while Neural Honeytrace remains effective with the minimal WSR of
Hyperparameters. Subsequently, we evaluate how the hyperparameters will affect the performance of Neural Honeytrace. Specifically, we consider the query sample size for model extraction attackers, and the three hyperparameters used in Neural Honeytrace.
As illustrated in the first column in Fig. 7, we perform different attacks with KnockoffNet and different sample sizes on the target model trained on CIFAR-10. As the sample size increases from to , the extraction accuracy of the stolen model slightly increases, because larger sample sizes help attackers gain more information about the feature space of the target model. At the same time, the watermark success rate remains stable under different sample sizes, which indicates that attackers cannot bypass Neural Honeytrace by adjusting the number of queries.
The other columns in Fig. 7 show the effectiveness of the three hyperparameters, , of Neural Honeytrace. These hyperparameters are used to balance the model availability and the watermark success rate. According to Eq. 7, Eq. 9, and Eq. 10, intuitively, larger and smaller leads to stronger watermarks but lower protected accuracy, which can also be observed in Fig. 7. Therefore, given a test dataset and the acceptable maximum drop in accuracy, model owners can identify appropriate hyperparameters.
Watermark detection and removal. In Sec. 5.2, we evaluate existing watermarking methods against adaptive model extraction attacks. However, previous research [9] suggests that attackers may employ test-time backdoor defense strategies to prevent triggerable watermarks from being activated. Therefore, we additionally evaluated Neural Honeytrace against different backdoor defenses.
For backdoor detection, we use BTI [34], a backdoor model detection method based on trigger inversion. We conduct multiple model extraction attacks on the target model trained on CIFAR-10, obtaining stolen models with Neural Honeytrace and models without any defense. We then calculate the average detection heatmap of BTI on these test models. As shown in Fig. 8, the detection heatmaps exhibit similar trends across class pairs, with no anomalous small values (which would suggest potential backdoors) pointing to the target class (class 9) in Neural Honeytrace. This indicates that Neural Honeytrace is robust against trigger-inversion-based backdoor detection.
For backdoor removal, we use CLP [35], a data-free backdoor removal method based on channel Lipschitzness. Specifically, we first perform model extraction attacks on the target model trained on CIFAR-10 and get the stolen model. Then we perform model pruning on the stolen model and record the watermark success rate. As shown in Fig. 9, as the pruning strength increases, the watermark success rate of Neural Honeytrace remains stable. The watermarking success rate does not decrease even if the cleaning accuracy of the stolen model decreases significantly. This indicates that the watermarks injected by Neural Honeytrace are strongly coupled to the original task and are difficult to remove.
| Method | Naive | Oracle | ||||
| Acc | Fidelity | WSR | Acc | Fidelity | WSR | |
| EWE | 88.71% | 88.18% | 39.90% | 86.02% | 87.95% | 4.40% |
| Composite | 85.47% | 87.41% | 43.80% | 85.25% | 87.63% | 10.40% |
| MEA-Defender | 88.15% | 88.63% | 61.80% | 86.10% | 87.95% | 10.80% |
| Neural H/T | 83.23% | 81.16% | 65.00% | 64.19% | 65.28% | 11.80% |
Oracle Attack. We also evaluate Neural Honeytrace against a highly capable attacker with full knowledge of the defense mechanism. Specifically, this attacker possesses a set of samples containing both the ground-truth predictions made by the target model and the corresponding modified predictions generated by watermarking strategies. Using these prediction pairs, the attacker trains a label recovery network following the method in [10] and subsequently trains a surrogate model using the recovered predictions.
As shown in Tab.IX, the watermark success rate of triggerable watermarking methods decreases significantly under oracle attacks compared to Naive Attack. However, the extraction accuracy and fidelity achieved by oracle attacks against Neural Honeytrace are substantially lower than those against previous methods. Furthermore, compared to perturbation-based defenses[11], Neural Honeytrace demonstrates even lower extraction accuracy and fidelity, indicating that the stolen model fails to accurately reconstruct the functionality of the target model. Therefore, Neural Honeytrace is more robust than existing methods when defending against oracle attacks. More importantly, model owners can easily change the watermarking details since Neural Honeytrace supports plug-and-play implementation, making it more difficult for attackers to access the defense information.
6 Related Work
6.1 Model Extraction Attack
Existing research reveals that machine learning models have different security risks throughout their lifecycle, threatening the availability [36, 37], integrity [38, 39], and confidentiality [40, 41]. Among these malicious attacks, model extraction attacks aim to breach the confidentiality of closed-source models for two goals: (1) rebuilding the functionality of the target model and use it without payment [1], or (2) conducting black-box attacks on the target model via surrogate models [42]. In this paper, we focus on model extraction attacks that attempt to steal functionality.
Naive Attacks. Previous work have explored using different query strategies to perform model extraction attacks. For example, KnockoffNet [4] and ActiveThief [43] select representative natural samples for querying. Subsequent research found that samples close to the decision boundary contain more parameter information, motivating utilizing synthetic (adversarial) samples to perform attacks. In practice, JBDA-TR [5] utilized the feedback of the target model to guide the synthesis process, while FeatureFool [44] used different adversarial attacks to generate query samples. With the development of data-free distillation techniques [45], some methods utilize generative models to generate query samples (e.g., MAZE [46] and MEGEX [47].)
Adaptive Attacks. To mitigate the threat of model extraction attacks, various defense strategies have been proposed, including detection-based defenses [5, 13] and perturbation-based defenses [48]. In response, several adaptive model extraction attacks have been developed to bypass these defenses. The S4L Attack [7] combines cross-entropy loss with a semi-supervised loss to extract more information with a limited number of queries. The Smoothing Attack [9] augments each sample multiple times and averages the predictions to train the model. D-DAE [10] trains both a defense detection model and a label recovery model to detect and bypass potential defenses. The p-Bayes Attack [11] utilizes independent and neighborhood sampling for real label estimation.
6.2 Model Watermarking
White-box watermarking. White-box watermarking methods aim at defending direct model stealing. For example, Uchida et al. [49] and Adi et al. [50] both embed watermarks in the parameter space. However, these methods require a white-box access to the stolen model, making them less practicable.
Black-box watermarking. In black-box conditions, defenders can only query the suspicious model through certain interfaces. Namba et al. [51] use a set of sample-label pairs to embed a backdoor-like watermark in the target model. The subsequent method SSL-WM [52] migrated this method to protect pretrained models by injecting task-agnostic backdoors. Nevertheless, these methods are ineffective against model extraction attacks, because the watermarks fail to transmit from protected models to stolen models.
Szyller et al. [17] proposed a simple strategy against model extraction attacks by randomly mislabeling some input queries. Other methods attempted to enhance the success rate of watermark transmission [16, 19, 20]. For example, EWE [16] links watermarking learning to main task learning closely by adding additional regular terms. On the basis of EWE, MEA-Defender [20] adopted the composite backdoor [18] as watermarks, further enhancing the success rate of watermark transmission. However, there is still no effective theoretical framework for watermarking against model extraction attacks.
7 Conclusion, Limitations, and Future Work
In this paper, we propose Neural Honeytrace, a robust plug-and-play watermarking framework. We model watermark transmission problem using information theory. Based on the analysis, we propose two watermarking strategies: training-free watermark embedding and multi-step watermark transmission, to achieve training-free and robust watermarking. Experimental results show that Neural Honeytrace is significantly more robust and provides flexible.
Although Neural Honeytrace mostly maintains the availability of the protected model, it introduces slight performance decrement, which poses the need for distinguishing benign and malicious queries. We adopts a simple algorithm, which may fail when attackers have access to the training dataset. One potential solution is to dynamically adjust the parameters for suspicious and trusted users based on their historical behavior.
The most relevant future work is to implement Neural Honeytrace on generative models such as Stable-Diffusion. Although we mainly evaluated Neural Honeytrace on classification tasks following existing methods, our watermark transmission model suggests that the larger output space of generative models may be able to effectively transmit more watermark information in one query.
Appendix A Additional Experimental Results

### v4

@@ 4.2 Multi-step Watermark Transmission
Training-free watermark embedding enables plug-and-play watermarking, however, the channel capacity is not always ideal for watermark transmission. Therefore, we propose the multi-step watermark transmission strategy to enhance the robustness of Neural Honeytrace.
@@
where is the reference logits of the target class, is used to balance model availability and watermark success rate, Bernoulli randomly samples with probability to decide whether to flip the label, and is a small random value to maintain randomness. Intuitively, for , the input sample will be labeled as the target class with a high probability. And by controlling , defenders can determine the flipping ratio for samples with . Eq. 7 links the predicted labels to the watermark similarities.
Robustness analysis. Label-flipping-based watermark information transmission embeds similarity scores in the label distribution of a set of samples with similar similarity scores. Consider the queries of these samples as a one-time message transmission, the channel capacity satisfies . And in practice, is usually a large number because attackers will utilize many samples to query the target model. Consequently, enables successfully watermark transmission.
5 Experiments
5.1 Experiment Setup
Datasets. We use four different image classification datasets to train the target model: CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), Caltech-256 (Griffin et al., 2007), and CUB-200 (Wah et al., 2011).
We use another two datasets as surrogate datasets used by model extractions attackers, TinyImageNet-200 (Mnmoustafa, 2017) for querying target models trained on CIFAR-10 and CIFAR-100, ImageNet-1K (Deng et al., 2009) for querying target models trained on Caltech-256 and CUB-200.
Models. We use two model architectures to train target models: VGG16-BN (Simonyan and Zisserman, 2015) for CIFAR-10 and CIFAR-100, and ResNet50 (He et al., 2016) for Caltech-256 and CUB-200. Following previous model extraction defenses (Orekondy et al., 2020; Tang et al., 2024), the same architectures are used for surrogate models.
Metrics. We use three metrics to evaluate the effectiveness of different watermarking methods: (1) Protected Accuracy indicates the accuracy of the protected model on clean samples, i.e., clean accuracy. (2) Extracted Accuracy (E-Acc) indicates the clean accuracy of the stolen model reconstructed by attackers. (3) Watermark Success Rate (WSR) measures the proportion of successfully activated watermarks. Specifically, it is the ratio of samples with added triggers that are classified into the target classes by the watermarked model but classified into different classes by non-protected models.
Model extraction attack methods. We consider two different query strategies, KnockoffNet (Orekondy et al., 2019) and JBDA-TR (Juuti et al., 2019), as basic model extraction attack methods. Additionally, we consider the following adaptive attack methods on the basis of KnockoffNet and JBDA-TR: S4L Attack (Jagielski et al., 2020), Smoothing Attack (Lukas et al., 2022), D-DAE (Chen et al., 2023), p-Bayes Attack (Tang et al., 2024), and Top-1 Attack (Only hard-labels are used for training surrogate models).
Watermarking strategies. We compare Neural Honeytrace with the following watermarking strategies in experiments: (1) DAWN (Szyller et al., 2021): Defenders randomly flip the predicted label of a subset of input queries and record these sample-label pairs as watermarks. In the default configuration, we assume that of all queries are executed by model extraction attackers. (2) EWE (Jia et al., 2021): Defenders utilize the Soft Nearest Neighbor Loss (SNNL) to minimize the distance between watermark features and natural features. (3) Composite Backdoor (Lin et al., 2020): Spliced in-distribution samples are used as watermarks (triggers) to increase the watermark success rate. (4) MEA-Defender (Lv et al., 2024b): Defenders introduce the utility loss, the watermarking loss, and the evasion loss to balance the model availability and watermark success rate.
@@ 5.1 Experiment Setup
| Protected Accuracy | 81.95% | 82.40% | 82.41% | 82.67% | 80.13% | ||||||
5.2 Experimental Results
We begin by comparing Neural Honeytrace with four baseline methods. For each defense, we report both the average and minimum Watermark Success Rate (WSR). In practice, watermarks must remain effective even at the minimum WSR to ensure security. We make the following observations based on the results:
Existing watermarking methods are vulnerable to adaptive attacks. As shown in Tab. 2 and Tab. 3, five adaptive attack strategies can weaken the watermark to varying degrees. Overall, the Smoothing Attack and Top-1 Attack have the most significant impact on WSR when KnockoffNet and JBDA-TR are used as query strategies, respectively.
MEA-Defender outperforms other baseline defenses. Among four baseline methods, although DAWN achieves the highest average WSR across four datasets, it relies on the assumption that model extraction queries occur in a high percentage of all queries, which is not satisfied in most cases. Compared with EWE and Composite Backdoor, MEA-Defender shows higher a average WSR and better robustness against most adaptive attacks. However, the minimum WSR for MEA-Defender indicates that it will still fail under certain adaptive attacks.
Neural Honeytrace outperforms existing defenses. Compared to existing watermarking strategies, Neural Honeytrace achieves better transferability across datasets. More importantly, Neural Honeytrace’s minimal WSR ensures its effectiveness even under worst-case scenarios, with copyright declaration overhead less than of existing methods.
| Acc | WSR | Acc | WSR | Acc | WSR | Acc | WSR | Acc | WSR | Acc | WSR | ||
| Attack | V-16R-18 | V-16R-50 | R-18V-16 | |||
| Acc | WSR | Acc | WSR | Acc | WSR | |
| Naive | 54.04% | 47.60% | 56.44% | 55.80% | 65.67% | 77.80% |
| S4L | 52.50% | 55.00% | 53.46% | 61.40% | 62.29% | 80.80% |
| Smoothing | 53.74% | 58.20% | 56.68% | 59.20% | 65.03% | 73.40% |
| DDAE | 41.93% | 54.80% | 44.76% | 56.20% | 65.59% | 68.20% |
| P-Bayes | 52.66% | 43.80% | 55.18% | 49.40% | 64.76% | 70.80% |
| Top-1 | 46.10% | 45.00% | 50.03% | 51.80% | 55.31% | 65.60% |
5.3 Ablation Study
Different Watermark Triggers. In Tab. 4, we evaluate three different watermark triggers: white pixel blocks (used in EWE (Jia et al., 2021)), a semantic object (e.g., a specific copyright logo), and a composite trigger (the default configuration). As shown in Tab. 4, the composite trigger achieves the highest average WSR among the three triggers, which aligns with the analysis in Sec. 4.1. Compared to the semantic object trigger, white pixel blocks yield a higher average WSR due to their simpler features.
Different Query Datasets. In Tab. 4, we compare the performance of Neural Honeytrace when attackers use CIFAR-10, CIFAR-100, and TinyImageNet as surrogate datasets, respectively. The experimental results show that for out-of-distribution surrogate datasets (CIFAR-100 and TinyImageNet), Neural Honeytrace maintains high WSR. When using the same training dataset as the surrogate (CIFAR-10), the WSR decreases but remains sufficiently high for ownership claims within queries.
Different Model Architectures. Tab. 5 lists the performance of Neural Honeytrace on different teacher-student model pairs. Combining Tab. 5 and Tab. 2, it can be observed that when the model structures are different, the stolen accuracy will slightly decrease, while Neural Honeytrace remains effective with the minimal WSR of
Hyperparameters. We evaluate different query sample size for model extraction attackers, and the three hyperparameters used in Neural Honeytrace.
As illustrated in the first column in Fig. 5, as the sample size increases from to , the extraction accuracy of the stolen model slightly increases, while the WSR remains stable under different sample sizes.
The other columns in Fig. 5 show the effectiveness of the three hyperparameters, , of Neural Honeytrace. According to Eq. 4, Eq. 6, and Eq. 7, intuitively, larger and smaller leads to stronger watermarks but lower protected accuracy, which can also be observed in Fig. 5.
| Attack | Setting | CIFAR-10 | CIFAR-100 | ||
| Acc | WSR | Acc | WSR | ||
| DFME | Soft-label | 66.53% | 53.40% | 38.29% | 49.80% |
| DFMS | Soft-label | 71.81% | 62.00% | 40.78% | 57.40% |
| Hard-label | 68.69% | 68.20% | 36.45% | 66.60% | |
| DisGUIDE | Soft-label | 76.48% | 57.80% | 45.58% | 64.20% |
| Hard-label | 73.10% | 64.20% | 40.44% | 70.80% | |
Data-free Model Extraction Attacks. Data-free model extraction attacks generate synthetic data to replicate a victim model’s behavior without access to real training data. Tab. 6 lists the performance of Neural Honeytrace against three different data-free MEAs, DFME (Truong et al., 2021), DFMS (Sanyal et al., 2022), and DisGUIDE (Rosenthal et al., 2023). The experimental results show that Neural Honeytrace is effective for defending all three kinds of data-free MEAs, with the average WSR of . This is because Neural Honeytrace does not rely on prior knowledge about the query dataset and is data-irrelevant.
Watermark detection and removal. Attackers may employ backdoor defense strategies to prevent triggerable watermarks. For backdoor detection, we use BTI (Tao et al., 2022), a backdoor model detection method based on trigger inversion. As shown in Fig. 6, the detection heatmaps exhibit similar trends across class pairs, with no anomalous small values pointing to the target class (class 9). For backdoor removal, we use CLP (Zheng et al., 2022), a data-free backdoor removal method based on channel Lipschitzness. As shown in Fig. 7, as the pruning strength increases, the WSR of Neural Honeytrace remains stable.
6 Related Work
6.1 Model Extraction Attack
Model extraction attacks aim to breach the confidentiality of closed-source models for two goals: (1) rebuilding the functionality of the target model and use it without payment (Tramèr et al., 2016), or (2) conducting black-box attacks on the target model via surrogate models (Liu et al., 2022). In this paper, we focus on model extraction attacks that attempt to steal functionality.
Naive Attacks. Previous work have explored using different query strategies to perform model extraction attacks. For example, KnockoffNet (Orekondy et al., 2019) and ActiveThief (Pal et al., 2020) select representative natural samples for querying. Subsequent research found that samples close to the decision boundary contain more parameter information, motivating utilizing synthetic (adversarial) samples to perform attacks. In practice, JBDA-TR (Juuti et al., 2019) utilized the feedback of the target model to guide the synthesis process, while FeatureFool (Yu et al., 2020) used different adversarial attacks to generate query samples. With the development of data-free distillation techniques (Yu et al., 2023), some methods utilize generative models to generate query samples (e.g., MAZE (Kariyappa et al., 2021) and MEGEX (Miura et al., 2024).)
Adaptive Attacks. Several adaptive model extraction attacks have been developed to bypass potential defenses. The S4L Attack (Jagielski et al., 2020) combines cross-entropy loss with a semi-supervised loss to extract more information with a limited number of queries. The Smoothing Attack (Lukas et al., 2022) augments each sample multiple times and averages the predictions to train the model. D-DAE (Chen et al., 2023) trains both a defense detection model and a label recovery model to detect and bypass potential defenses. The p-Bayes Attack (Tang et al., 2024) utilizes neighborhood sampling for real label estimation.
6.2 Black-box Model Watermarking
In black-box conditions, defenders can only query the suspicious model through certain interfaces. Namba et al. (Namba and Sakuma, 2019) use a set of sample-label pairs to embed a backdoor-like watermark in the target model. The subsequent method SSL-WM (Lv et al., 2024a) migrated this method to protect pretrained models by injecting task-agnostic backdoors. Nevertheless, these methods are ineffective against model extraction attacks, because the watermarks fail to transmit to stolen models.
Szyller et al. (Szyller et al., 2021) proposed a simple strategy against model extraction attacks by randomly mislabeling some input queries. Other methods attempted to enhance the success rate of watermark transmission (Jia et al., 2021; Cong et al., 2022; Lv et al., 2024b). For example, EWE (Jia et al., 2021) links watermarking learning to main task learning closely by adding additional regular terms. On the basis of EWE, MEA-Defender (Lv et al., 2024b) adopted the composite backdoor (Lin et al., 2020) as watermarks, further enhancing the success rate of watermark transmission.
7 Conclusion
In this paper, we propose Neural Honeytrace, a robust plug-and-play watermarking framework. We model watermark transmission problem using information theory. Based on the analysis, we propose two watermarking strategies: training-free watermark embedding and multi-step watermark transmission, to achieve training-free and robust watermarking. Experimental results show that Neural Honeytrace is significantly more robust and provides flexible solutions.
Appendix A Additional Experimental Results
A.1 Performance Comparison
| Query | Attack | DAWN | EWE | Composite | MEA-Defender | Ours | |||||
| E-Acc | WSR | E-Acc | WSR | E-Acc | WSR | E-Acc | WSR | E-Acc | WSR | ||
@@ A.1 Performance Comparison
Tab. 7 and Tab. 8 list the experimental results on CIFAR-100 and Caltech-256 datasets. Fig. 8 provides a visualization of the sample sizes required by different watermarking strategies for ownership claim. Considering both the average and worst case scenarios, Neural Honeytrace requires fewer queries compared to previous methods. For example, in the worst-case scenario, MEA-Defender and Neural Honeytrace require 193,252 and 1,857 queries respectively to achieve copyright declaration.